A body moving with a constant acceleration of A starting from 1m/s to 7m/s &it traveled a distance of S,what is the velocity at the middle of the trip?

Grade:12

1 Answers

Menka Malguri
12 years ago

u=1m/s

v=7m/s

distance travelled=S

v2-u2=2as

49-1=2aS

2aS=48

aS=24..................(1)

a=24/S

Distance travelled in middle of trip=12/a

Let the speed at this distance be V.

Then,

V2-u2=2aS

V2-1=2(24/S)(12/a).............(from(1))

V2=25

V=5m/s.